# A Systematic Search of Distant Superclusters with the Subaru Hyper Suprime-Cam Survey

### v1

@@ Abstract
[…]
@@ 1 Introduction
[…]
@@ 2.1 CAMIRA cluster catalog
The CAMIRA catalog is built by applying the CAMIRA algorithm (Oguri, 2014) to the internal release of HSC survey data “S21A” (with the latest bright star masks), with data obtained between April 2014 and January 2021.111The PDR3 is based on data obtained up to January 2020. In short, the CAMIRA algorithm utilizes the calibrated stellar population synthesis (SPS) model of Bruzual & Charlot (2003) to estimate the likelihood of a galaxy being a red-sequence galaxy and derived a “number parameter” based on it for each galaxy. Two-dimensional richness maps are then constructed by convolving a stellar mass filter and a spatial filter with the number parameter of each galaxy at a given redshift slice. For each peak in the richness maps, the CAMIRA cluster photo-$z$ $z_{CAM}$ is obtained by maximizing the likelihood (see Eqn. 15 of Oguri 2014) and iteratively identifying the Brightest Cluster Galaxy (hereafter BCG) position until convergence. Oguri et al. (2018) have shown that the redshifts are highly accurate, with a bias of $|\delta_{z}|\approx 0.005(1+z)$ and a scatter of $\sigma_{z}\approx 0.01(1+z)$ throughout the redshift range of $z=0.1-1.0$. For full details of the algorithm and performance of CAMIRA, please refer to Oguri (2014) and Oguri et al. (2018).
@@
(where $\sigma_{z}\approx 0.01$ is the scatter), we adopt $z_{BCG}$ as the cluster redshift.
Although the CAMIRA cluster catalog contains cluster candidates with photo-$z$ up to $z\approx 1.4$, we focus on the redshift range between $0.5$ and $1.0$ in this work, as the richness of $z>1$ CAMIRA clusters is not corrected for incompleteness (Oguri, 2014; Oguri et al., 2018) and therefore would be underestimated. The cluster sample selected in this redshift range contains 14,694 clusters over the HSC wide layer. Figure 1 shows the comoving number density of the CAMIRA clusters as a function of redshift. Compared to the halo number density predicted by the halo mass function of Tinker et al. (2008), the richness cuts of $N_{mem}\geq 10.0$ and $N_{mem}\geq 15.0$ correspond to halo masses $M_{200m}\approx 6.5\times 10^{13}\,h^{-1}\,M_{\odot}$ and $\approx 1.0\times 10^{14}\,h^{-1}\,M_{\odot}$, respectively.
@@ 2.2 DESI EDR
The cluster photo-$z$ validation is done as follows. For each cluster in our sample, we count the number of targets from DESI EDR within a cylindrical volume. The cylinder has a radius of $1h^{-1}\,$pMpc from the center of the cluster and a line-of-sight distance of ±$\pm 2500\,$km/s from the cluster photo-$z$ $z_{CAM}$. We regard the cluster photo-$z$ to be reliable if at least 5 targets lie within the cylinder. This way, photo-$z$ of 520 clusters are validated.
@@ 2.3 DEmP
We first obtain the DEmP photo-$z$ for each CAMIRA member galaxy by cross-matching the DEmP catalog and CAMIRA cluster member galaxy catalog with a separation less than $0.08^{\prime\prime}$ and select the nearest primary source. The photoz_mode is adopted as the point estimator of the redshift. For each CAMIRA cluster in our sample, we construct its photo-$z$ cumulative distribution function (CDF) after removing member galaxies with photoz_mode_risk $\geq 0.1$ and membership probability $w<w_{thres}$. We consider an interval centered at the CAMIRA photo-$z$, $[z_{l},z_{u}]$, where $z_{u}=z_{CAM}+\sigma_{z}(1+z_{CAM})$, $z_{l}=z_{CAM}-\sigma_{z}(1+z_{CAM})$, and $\sigma_{z}=0.01$. The cluster photo-$z$ is considered to be consistent with the photo-$z$ of the member galaxies if the following criteria hold: (1) $CDF(z_{u})-CDF(z_{l})>0.5$, and (2) Number of member galaxies within the interval $[z_{l},z_{u}]$ greater than or equal to $N_{min}$. The parameters $w_{thres}$ and $N_{min}$ are empirically determined; we adopt $w_{thres}=0.5$ and $N_{min}=10$ here.
@@
Figure 2: Demonstration of comparing CAMIRA cluster photometric redshift with DEmP photometric redshift of member galaxies. The blue curve shows the normalized cumulative distribution function (CDF) of member galaxies’ photo-$z$ estimated by DEmP. The black vertical curve indicates the cluster photo-$z$. The gray dashed line and the gray dash-dotted line indicate the value of CDF at the lower and upper bound of the redshift interval $[z_{l},z_{u}]$ (two vertical gray lines) considered, respectively. According to criteria described in Section 2.3, the cluster photo-$z$ is inconsistent with member galaxies’ photo-$z$ in the top and middle panels, while the bottom panel demonstrates a consistent case.
The upper panel of Figure 3 summarizes the result of our photo-$z$ examination using the three different approaches described in this section: consistent BCG spectroscopic redshift, DESI confirmation, and member galaxy DEmP photo-$z$ check. We refer to the subset of clusters that satisfy one or more of the confirmation tests as having confident redshifts. These clusters constitute one-third of the clusters selected from redshift $z=0.5-1.0$. The lower panel of Figure 3 shows the fraction of clusters that do not have a confident redshift as a function of the CAMIRA photo-$z$. This curve will later be utilized to perturb the redshift of the halo light cone to mimic the CAMIRA photo-$z$ uncertainty in Section 4.1.
@@ 3.1 N-body simulation
By definition, our supercluster candidates must undergo gravitational collapse eventually. We therefore rely on $N$-body simulations that are run well beyond $z=0$ (or $a=1$, where $a$ is the cosmic expansion factor) to find the optimal linking lengths for our FoF algorithm. We first generate the unigrid initial condition at redshift $z=100.0$ by MUSIC (MUlti-Scale-Initial-Conditions; Hahn & Abel 2011) with the transfer function provided by CAMB (Code for Anisotropies in the Microwave Background; Lewis et al. 2000). The initial condition is then evolved by GADGET-2 (Galaxies with dark matter and gas interact; Springel 2005) up to scaling factor $a=15.0$ when the age of the Universe is approximately 57.4 Gyr. Dark matter halos are extracted from each snapshot by the halo finder ROCKSTAR (Robust Overdensity Calculation using K-Space Topologically Adaptive Refinement; Behroozi et al. 2013). We have verified that the halo mass function output from ROCKSTAR is consistent with the analytical fitting formula of Tinker et al. (2008).
@@ 3.2 Construction of mock catalogs
Inspired by Lin et al. (in prep), we construct the mock catalog by assigning a richness to each halo in a Monte Carlo fashion. In practice, for a given ROCKSTAR halo catalog at redshift $z_{i}$, we consider halos with mass $M_{200m}\geq 2.0\times 10^{13}\,h^{-1}\,M_{\odot}$ and designate the richness $N_{mem}$ to each halo by drawing from the richness-halo mass relation $P(lnN|M=M_{200m},z=z_{i})$. The lower limit of the halo mass considered is selected such that less massive halos have a negligible probability of receiving richness $N_{mem}>10.0$, the lower limit of the CAMIRA catalog. After assigning richness to each halo, we apply the richness cut $N_{mem}\geq 10.0$ to construct mock catalogs. By this approach, we create ten Monte Carlo realizations for each of the six $N$-body simulations; therefore we have 60 different realizations in total.
@@
3.3 Optimzation of the linking length
@@
In this work, we empirically determine the linking length from a systematic comparison of the FoF groups (denoted $S_{FoF}(l)$) obtained by applying different linking lengths to the ground-truth superclusters (denoted $S_{TSC}$) in our $N$-body simulations. The ground-truth superclusters are the groups of halos in mock catalogs that will eventually collapse in the future (by $a=15.0$). In practice, for a mock catalog at a given redshift $z_{i}$, we trace the descendants following the merger trees of the halos. If a group of halos merges into one halo before or at $a=15.0$, that is, sharing the same parent ID or ID, then it is identified as a bona fide supercluster. On the other hand, we study the impact of the linking length by looking at the sets of FoF groups obtained with each choice of linking length. We require that both FoF groups and ground-truth superclusters have at least two member halos. For a given linking length $l$, we evaluate the performance of the FoF algorithm by two metrics, namely purity $P$ and completeness $C$:
$P(l\,;z_{i}\,,N_{cut})=\frac{\sum M_{200m}(S_{FoF}(l)\cap S_{TSC})}{\sum M_{200m}(S_{FoF}(l))}$ (2)
$C(l\,;z_{i}\,,N_{cut})=\frac{\sum M_{200m}(S_{FoF}(l)\cap S_{TSC})}{\sum M_{200m}(S_{TSC})}$ (3)
where $z_{i}$ is the redshift of the snapshot, $N_{cut}$ (say, 10.0) is the richness cut. The denominators of purity and completeness represent the total FoF groups halo mass and the total mass of the ground-truth supercluster member halos, respectively. The numerator [^2] We note that the $\cap$ symbol used in Equations 2 & 3 are not exactly the way it is used in mathematics. Please see the following text for a more precise description. is calculated by summing $M_{200m}$ of FoF groups that belong to one of the ground-truth superclusters. We note that it is possible that a ground-truth supercluster from $S_{TSC}$ is fragmented into more than two FoF groups in $S_{FoF}(l)$. In this case, we only sum the halo mass of the most massive FoF group in the numerator. This way, the purity $P$ quantifies the fraction of the total halo mass of $S_{FoF}(l)$ in ground-truth superclusters for a given $l$. The completeness $C$ quantifies the fraction of the total halo mass of $S_{TSC}$ recovered for a given $l$.

### v2

@@ Abstract
Superclusters, encompassing environments across a wide range of overdensities, can be regarded as unique laboratories for studying galaxy evolution. Although numerous supercluster catalogs have been published, none of them goes beyond redshift $z=0.7$. In this work, we adopt a physically motivated supercluster definition, requiring that superclusters should eventually collapse even in the presence of dark energy. Applying a friends-of-friends (FoF) algorithm to the CAMIRA cluster sample constructed using the Subaru Hyper Suprime-Cam survey data, we have conducted the first systematic search for superclusters at $z=0.5-1.0$ and identified 673 supercluster candidates over an area of 1027 deg2. The FoF algorithm is calibrated by evolving $N$-body simulations to the far future to ensure high purity. We found that these high-$z$ superclusters are mainly composed of $2-4$ clusters, suggesting the limit of gravitationally bound structures in the younger Universe. In addition, we studied the properties of the clusters and brightest cluster galaxies (BCGs) residing in different large-scale environments. We found that clusters associated with superclusters are typically richer, but no apparent dependence of the BCG properties on large-scale structures is found. We also compared the abundance of observed superclusters with mock superclusters extracted from halo light cones, finding that photometric redshift uncertainty is a limiting factor in the performance of superclusters detection.
@@ 1 Introduction
The Hyper Suprime-Cam Subaru Strategic Program (HSC-SSP, hereafter the HSC survey) is a large optical imaging survey primarily designed to probe the dark matter distribution via weak lensing and to unveil the evolution of galaxies (Aihara et al., 2018a, b). Using the Hyper Suprime-Cam (Miyazaki et al., 2012, 2018; Komiyama et al., 2018; Furusawa et al., 2018) installed at the $8.2$m Subaru telescope in Hawai’i, the 330-night survey started in 2014 and finished the observations in 2022. It consists of three layers: wide, deep, and ultra-deep, characterized by different survey areas and depths in five broad-band filters ($grizy$; Kawanomoto et al. 2018). The HSC survey data are processed through the hscPipe pipeline (Bosch et al., 2018; Huang et al., 2018), which originates from the Legacy Survey of Space and Time (LSST) pipeline (Jurić et al., 2017; Ivezić et al., 2019). To conduct astrometric and photometric calibration, the HSC survey has utilized Pan-STARRS1 data (Tonry et al., 2012; Schlafly et al., 2012; Magnier et al., 2013). In the latest public data release (PDR3; Aihara et al. 2022), the HSC-wide layer has reached a limiting magnitude of $i_{\rm AB}\approx 26$ ($5\sigma$ for point sources within a $2^{\prime\prime}$ aperture) over $\approx 670$ deg2 observed in $grizy$ filters.
With the extraordinary depth and area of the HSC survey, we have conducted a systematic search for superclusters at redshift $z=0.5-1.0$ over a sky area of $\approx 1027$ deg2 based on the latest internal data release (see Section 2.1). This is the first systematic search of distant superclusters that extends to such a high redshift over a thousand deg2. Our final supercluster catalog consists of 673 supercluster candidates with multiplicity between $2-5$.
@@ 2.1 CAMIRA cluster catalog
The CAMIRA catalog is built by applying the CAMIRA algorithm (Oguri, 2014) to the internal release of HSC survey data “S21A” (with the latest bright star masks), with data obtained between April 2014 and January 2021.111The PDR3 is based on data obtained up to January 2020. In short, the CAMIRA algorithm utilizes the calibrated stellar population synthesis (SPS) model of Bruzual & Charlot (2003) to estimate the likelihood of a galaxy being a red-sequence galaxy and derived a “number parameter” based on it for each galaxy. Two-dimensional richness maps are then constructed by convolving a stellar mass filter $F_{M}$ and a spatial filter $F_{R}$ with the number parameter of each galaxy at a given redshift slice. For each peak in the richness maps, the CAMIRA cluster photo-$z$ $z_{CAM}$ is obtained by maximizing the likelihood (see Eqn. 15 of Oguri 2014) and iteratively identifying the Brightest Cluster Galaxy (hereafter BCG) position until convergence. Oguri et al. (2018) have shown that the redshifts are highly accurate, with a bias of $|\delta_{z}|\approx 0.005(1+z)$ and a scatter of $\sigma_{z}\approx 0.01(1+z)$ throughout the redshift range of $z=0.1-1.0$. For full details of the algorithm and performance of CAMIRA, please refer to Oguri (2014) and Oguri et al. (2018).
@@
(where $\sigma_{z}\approx 0.01$ is the scatter), we adopt $z_{BCG}$ as the cluster redshift. In Appendix C.2, we provide an analysis showing that the redshift of BCG could be representative of cluster redshift.
Although the CAMIRA cluster catalog contains cluster candidates with photo-$z$ up to $z\approx 1.4$, we focus on the redshift range between $0.5$ and $1.0$ in this work, as the richness of $z>1$ CAMIRA clusters is not corrected for incompleteness (Oguri, 2014; Oguri et al., 2018) and therefore would be underestimated. Figure 1 shows the comoving number density of the CAMIRA clusters as a function of redshift. Compared to the halo number density predicted by the halo mass function of Tinker et al. (2008), the richness cuts of $N_{mem}\geq 10.0$ and $N_{mem}\geq 15.0$ correspond to halo masses $M_{200m}\approx 6.5\times 10^{13}\,h^{-1}\,M_{\odot}$ and $\approx 1.0\times 10^{14}\,h^{-1}\,M_{\odot}$, respectively.
@@ 2.2 DESI EDR
The cluster photo-$z$ validation is done as follows. Considering that cluster photo-$z$ $z_{CAM}$ might be offset from the true value, we first search for DESI galaxy concentrations by calculating the $1\sigma$-clipped median $z_{0}$ of all DESI galaxies within a “long” cylinder of radius $1h^{-1}\,$pMpc and line-of-sight distance of ±$\pm\Delta z(1+z_{CAM})$. We then place a “shorter” cylinder with radius $1h^{-1}\,$pMpc and line-of-sight distance of $\Delta v=\pm 3000\,$km/s at the location of $1\sigma$-clipped median $z_{0}$ derived above. If we find at least $N_{min}$ DESI galaxies within the shorter cylinder of a given cluster, we then designate the cluster as confirmed and derive a new redshift $z_{cl,DESI}$ for it using the median spectroscopic redshift of the DESI galaxies that falls inside the shorter cylinder. We select $\Delta z=0.02$, $N_{min}=4$, and detail the choice of these parameters in Appendix C.1. This way, photo-$z$ of 883 clusters are validated.
@@ 2.3 DEmP
We first obtain the DEmP photo-$z$ for each CAMIRA member galaxy by cross-matching the DEmP catalog and CAMIRA cluster member galaxy catalog with a separation less than $0.08^{\prime\prime}$ and select the nearest primary source  [^2] Using object IDs to retrieve the photometric redshift of cluster member galaxies is a more straightforward approach. Unfortunately, information on object IDs is not provided in the CAMIRA cluster member galaxy catalog when the work is done.. The photoz_mode is adopted as the point estimator of the redshift. For each CAMIRA cluster in our sample, we construct its photo-$z$ cumulative distribution function (CDF) after removing member galaxies with photoz_mode_risk $\geq 0.1$ and membership probability $w<w_{thres}$. We consider an interval centered at the CAMIRA photo-$z$, $[z_{l},z_{u}]$, where $z_{u}=z_{CAM}+\Delta z(1+z_{CAM})$, $z_{l}=z_{CAM}-\Delta z(1+z_{CAM})$, and $\Delta z=0.02$. The cluster photo-$z$ is considered to be consistent with the photo-$z$ of the member galaxies if the following criteria hold: (1) $CDF(z_{u})-CDF(z_{l})>0.5$, and (2) Number of member galaxies within the interval $[z_{l},z_{u}]$ greater than or equal to $N_{min}$. The parameters $w_{thres}$ and $N_{min}$ are empirically determined; we adopt $w_{thres}=0.5$ and $N_{min}=10$ here. The justification on the choices of these hyperparameters is described in Appendix C.3. We also note that, due to the extraordinary depth of the HSC wide layer, the completeness is $>95\%$ in the redshift range $z=0.3-1.1$ for galaxies with stellar mass $\gtrapprox 3\times 10^{9}M_{\odot}$ (Lin et al., 2017). Therefore, we conduct the confirmation process above without considering the effect of completeness.
@@
Figure 2: Demonstration of comparing CAMIRA cluster photometric redshift with DEmP photometric redshift of member galaxies. The blue curve shows the normalized cumulative distribution function (CDF) of member galaxies’ photo-$z$ estimated by DEmP. The black vertical curve indicates the cluster photo-$z$. The gray dashed line and the gray dotted line indicate the value of CDF at the lower and upper bound of the redshift interval $[z_{l},z_{u}]$ (two vertical gray lines) considered, respectively. According to criteria described in Section 2.3, the cluster photo-$z$ is inconsistent with member galaxies’ photo-$z$ in the top and middle panels, while the bottom panel demonstrates a consistent case.
2.4 Summary of Photo-z examinations
@@
In the previous three subsections, we have used three different approaches to examine the photometric redshift accuracy of CAMIRA clusters. The upper panel of Figure 3 summarizes the results of these three examination process: confirmation with DESI, BCG spectroscopic redshift consistency examination, and member galaxy DEmP photo-$z$ check. The fraction of clusters validated by DESI EDR, BCG spec-$z$, and DEmP photo-$z$ are $6\%\,(883/14695)$, $17\%\,(2438/14695)$, and $56\%\,(8231/14695)$, respectively. The low fraction of DESI and BCG spec-z examination result from the limited area of available spectroscopic surveys. Although based on the same photometric catalog, the DEmP examination validates around half of the CAMIRA clusters in our sample. This is expected as photometric redshift of galaxies typically has a larger scatter ($\approx 0.025(1+z_{phot})$, see Nishizawa et al. 2020) than that of clusters. In addition, among the clusters validated by BCG spec-$z$, approximately $63\%$ of clusters are also validated by DEmP photo-$z$, which is close to our completeness estimation described in Appendix C.3.
We refer to the subset of clusters that satisfies one or more of the confirmation tests as having confident redshifts. These clusters constitute $64\%$ of the clusters at redshift $z=0.5-1.0$. When a cluster is validated by multiple confirmation tests described above, we use the redshift of the cluster following the order: $z_{cl,DESI}$ (Section 2.2), $z_{BCG}$ (Section 2.1), and $z_{CAM}$ (Section 2.3). Following this redshift priority, 14,695 clusters at $z=0.5-1.0$ from the HSC wide layer are selected as our sample. The lower panel of Figure 3 shows the fraction of clusters that do not have a confident redshift as a function of the CAMIRA photo-$z$. This curve will later be utilized to perturb the redshift of the halo light cone to mimic the CAMIRA photo-$z$ uncertainty in Section 4.1.
@@ 3.1 N-body simulation
By definition, our supercluster candidates must undergo gravitational collapse eventually. We therefore rely on $N$-body simulations that are run well beyond $z=0$ (or $a=1$, where $a$ is the cosmic expansion factor) to find the optimal linking lengths for our FoF algorithm. We first generate the unigrid initial condition at redshift $z=100.0$ by MUSIC (MUlti-Scale-Initial-Conditions; Hahn & Abel 2011) with the transfer function provided by CAMB (Code for Anisotropies in the Microwave Background; Lewis et al. 2000). The initial condition is then evolved by GADGET-2 (Galaxies with dark matter and gas interact; Springel 2005) up to scaling factor $a=15.0$  [^3] The optimized linking length is only reduced by approximately $0.1\,h^{-1}\,$cMpc if we set the final snapshot at $a=10.0$. Hence, the optimization process converges with respect to the end of simulation time. when the age of the Universe is approximately 57.4 Gyr. Dark matter halos are extracted from each snapshot by the halo finder ROCKSTAR (Robust Overdensity Calculation using K-Space Topologically Adaptive Refinement; Behroozi et al. 2013). We have verified that the halo mass function output from ROCKSTAR is consistent with the analytical fitting formula of Tinker et al. (2008).
@@ 3.2 Construction of mock catalogs
Inspired by Lin et al. (in prep), we construct the mock catalog by assigning a richness to each halo in a Monte Carlo fashion. In practice, for a given ROCKSTAR halo catalog at redshift $z_{i}$, we consider halos with mass $M_{200m}\geq 2.0\times 10^{13}\,h^{-1}\,M_{\odot}$ and designate the richness $N_{mem}$ to each halo by drawing from the richness-halo mass relation $P(lnN|M=M_{200m},z=z_{i})$ (Murata et al., 2019). The lower limit of the halo mass considered is selected such that less massive halos have a negligible probability of receiving richness $N_{mem}>10.0$, the lower limit of the CAMIRA catalog. After assigning richness to each halo, we apply the richness cut $N_{mem}\geq 10.0$ to construct mock catalogs. By this approach, we create ten Monte Carlo realizations for each of the six $N$-body simulations; therefore we have 60 different realizations in total.
@@
3.3 Optimization of the linking length
@@
In this work, we empirically determine the linking length from a systematic comparison of the FoF groups obtained by applying different linking lengths to the ground-truth superclusters in our $N$-body simulations. The ground-truth superclusters are the groups of halos in mock catalogs that will eventually collapse in the future (by $a=15.0$). In practice, for a mock catalog at a given redshift $z_{i}$, we trace the descendants following the merger trees of the halos. If a group of halos merges into one halo before or at $a=15.0$, that is, sharing the same parent ID or ID, then it is identified as a bona fide supercluster. We collectively called the set containing all ground-truth superclusters at a given redshift $S_{TSC}$. On the other hand, we study the impact of the linking length by looking at the sets of FoF groups obtained with each choice of linking lengths. For a given linking length $l$, we collectively call the set containing all FoF groups $S_{FoF}(l)$. We require that each FoF group in $S_{FoF}(l)$ and each ground-truth supercluster in $S_{TSC}$ have at least two member halos. The set $S_{TSC}$ is linking length independent, while set $S_{FoF}(l)$ depends on the chosen linking length.
For a given linking length $l$, we evaluate the performance of the FoF algorithm by two metrics [^4] Although the performance metrics here are weighted by halo mass, we also have tried optimizing the linking length based on unweighted metrics and found that our results are unbiased. The details are presented in Appendix B., namely purity $P$ and completeness $C$:
$P(l\,;z_{i}\,,N_{cut})=\frac{\sum M_{200m}(S_{FoF}(l)\land S_{TSC})}{\sum M_{200m}(S_{FoF}(l))}$ (2)
$C(l\,;z_{i}\,,N_{cut})=\frac{\sum M_{200m}(S_{FoF}(l)\land S_{TSC})}{\sum M_{200m}(S_{TSC})}$ (3)
where $z_{i}$ is the redshift of the snapshot, $N_{cut}$ (e.g., 10.0) is the richness cut. The denominators of purity and completeness represent the total FoF group halo mass and the total mass of the ground-truth supercluster member halos, respectively. For the numerator of purity and completeness, we describe the operator $\land$ in detail using the examples below. Consider a ground-truth supercluster composed of four member halos, say TSC=[CL1, CL2, CL3, CL4], the following 4 cases are possible:
  1. 1.
    A FoF group with exactly the same composition as the TSC is captured by the FoF algorithm (i.e., FoF1=[CL1, CL2, CL3, CL4]). In this case, we include the total halo mass of FoF1 in the numerator.
  2. 2.
    A FoF group that contains part of the TSC is found by the FoF algorithm, say FoF1=[CL1, CL2, CL3]. In this case, we add the total halo mass of FoF1 to the numerator.
  3. 3.
    The TSC is identified as two FoF groups, say FoF1=[CL1, CL2] and FoF2=[CL3, CL4]. In this case, we only consider the more massive FoF group of the two in the numerator, so that the fragmentation situation degrades both purity and completeness.
  4. 4.
    A FoF group could contain additional halo members other than halo members of TSC, say FoF1=[CL1, CL2, CL3, CL5]. In this case, FoF1 will not be added to the numerator so that the completeness does not increase monotonically.
Following the rules described above, each FoF group in the “intersection” set, $S_{FoF}(l)\land S_{TSC}$, is contained in a ground-truth supercluster from $S_{TSC}$. This way, the purity $P$ quantifies the fraction of the total halo mass of $S_{FoF}(l)$ in ground-truth superclusters for a given $l$. The completeness $C$ quantifies the fraction of the total halo mass of $S_{TSC}$ recovered for a given $l$.